### v1

# Are LLMs Really Not Knowledgable? Mining the Submerged Knowledge in LLMs’ Memory

Abstract
Large language models (LLMs) have shown promise as potential knowledge bases, yet they often struggle with question-answering tasks and are prone to hallucinations. While previous research attributes these issues to knowledge gaps in the model’s parameters, our investigation reveals a different phenomenon: LLMs often retain correct knowledge even when generating incorrect answers. Through analysis of model’s internal representations, we find that correct answers frequently appear among high-probability tokens despite not being selected as final outputs. Based on this observation, we introduce Hits@$k$, a new metric to assess knowledge retention independent of expression accuracy. Our extensive experiments demonstrate that LLMs store significantly more knowledge than their QA performance suggests. Building on these findings, we develop SkipUnsure, a method to improve answer accuracy by leveraging detected but unexpressed knowledge. Experiments on both open-domain and specific-domain datasets show consistent improvements, with accuracy gains of up to 11.8% on DBPedia and 6.3% on IMDB, without requiring model retraining.
1 Introduction
Large language models (LLMs; Touvron et al. 2023a; Chiang et al. 2023; Almazrouei et al. 2023; MosaicML 2023; Touvron et al. 2023b; OpenAI 2022; Google 2023) have emerged as potential alternatives to traditional knowledge bases, demonstrating capabilities in encoding and retrieving vast amounts of factual information through their parameters. The ability to accurately access and utilize this knowledge is crucial for reliable deployment of LLMs in real-world applications, from question answering to decision support systems. However, these models frequently produce incorrect answers or hallucinations in knowledge-intensive tasks, severely limiting their practical utility. Recent studies have explored multiple approaches to enhance LLMs’ knowledge utilization, including domain-specific fine-tuning Kumar et al. (2024), prompt engineering strategies Zhang et al. (2023a), and architectural modifications Zhong et al. (2023). These methods operate under the assumption that answer inaccuracies stem from insufficient knowledge in model parameters, leading to solutions focused on expanding model capacity or training data.
Our systematic investigation reveals fundamental limitations in this understanding of LLMs’ knowledge utilization. Analysis of model outputs demonstrates that even when generating incorrect answers, LLMs often maintain access to accurate information within their probability distributions over vocabulary tokens. In state capital queries, for instance, while models might output “Seattle” as Washington’s capital, they consistently assign high probability scores to the correct answer “Olympia”. This pattern persists across various knowledge domains and model architectures, indicating a systematic gap between knowledge storage and expression rather than simple knowledge absence.
To quantify this phenomenon, we propose Hits@$k$ to evaluate knowledge retention independent of answer accuracy. Extensive experiments across multiple datasets demonstrate the prevalence of this storage-expression gap. On DBpedia, LLaMA3-8b achieves only 17.2% standard accuracy (Hits@1) but reaches 57.9% for Hits@5, revealing substantially more stored knowledge than conventional metrics suggest. This disparity is particularly pronounced in domain-specific tasks and varies systematically with data popularity, offering insights into how LLMs organize and access their stored knowledge. Traditional evaluation methods, focusing solely on final outputs, significantly underestimate the knowledge actually encoded in model parameters.
Based on these insights, we develop SkipUnsure, an approach that improves answer accuracy by effectively utilizing knowledge present in model representations. Unlike existing methods requiring substantial computational resources or architectural changes, our approach identifies and filters uninformative responses while recovering high-probability knowledge from internal distributions. Experiments demonstrate consistent improvements across diverse scenarios, achieving accuracy gains of 11.8% on DBPedia and 6.3% on IMDB. These improvements are particularly pronounced for head entities where models show strong knowledge retention but poor expression.
This work makes several key contributions:
  • Identification and analysis of the systematic gap between knowledge storage and expression in LLMs;
  • Introduction of Hits@$k$ for quantifying model knowledge retention;
  • Comprehensive examination of factors influencing knowledge storage and expression patterns;
  • Development of SkipUnsure, demonstrating practical benefits of leveraging stored but unexpressed knowledge.
Figure 1: An example illustrating a scenario where a model possesses potentially correct memories yet fails to provide the correct answer.
2 Exploring Memory in LLMs
2.1 Knowledge Storage and Expression
Recent studies have explored using LLMs as knowledge bases, highlighting their potential to encode information within parameters through pre-training Petroni et al. (2019); Wang et al. (2020). While these models demonstrate impressive capabilities in question answering tasks, they often struggle with consistency and hallucination. Prior work frequently attributes such failures to knowledge gaps in the model’s parameters Sun et al. (2023); Li et al. (2024), suggesting that expanding model capacity or training data could address these issues.
Our investigation reveals that model failures may stem from expression issues rather than knowledge gaps. Through systematic analysis of model outputs, we find that LLMs often retain correct information in their parameters even when generating incorrect answers. As shown in Figure 1, when asked about Washington state’s capital, while the model outputs “Seattle”, it assigns a high probability score to the correct answer “Olympia”. Such cases indicate the need for a deeper understanding of how knowledge is stored and expressed in these models.
2.2 Analyzing Model’s Internal Knowledge
We investigate this phenomenon by examining the logits, which represent token probabilities produced during the model’s answer generation process. In LLMs, these logits reflect the model’s internal knowledge state before the final output selection. Our analysis of these distributions reveals a consistent pattern. Even when the model fails to output the correct answer, it often assigns significant probability scores to tokens representing the correct information. This observation persists across various question types and knowledge domains. The pattern is particularly evident in specialized domains, where models might respond with “unsure” while assigning high probabilities to correct technical terms. This suggests that traditional evaluation methods focusing solely on the model’s final output may substantially underestimate the knowledge actually stored in the model’s parameters.
2.3 The Hits@$k$ Metric
Building on these observations, we propose the Hits@$k$ metric to quantify the model’s knowledge retention:
$\text{Hits}@k=\frac{N^{k}_{correct}}{N}$ (1)
where $N^{k}_{correct}$ represents cases where the correct answer appears within the top-k logits. For large vocabulary models such as LLaMA3 with approximately 128,000 tokens, we find that a relatively small k value effectively captures stored knowledge while maintaining computational efficiency.
Experimental results in Figure 4 demonstrate the effectiveness of this metric in revealing stored knowledge. Using LLaMA3-8b on DBpedia, while Hits@1 is only 17.2%, Hits@5 reaches 57.9%, indicating substantially more stored knowledge than suggested by traditional metrics. This pattern holds across different domains and data types, suggesting a fundamental characteristic of how LLMs store and access information. These findings motivate a deeper examination of factors affecting knowledge storage and expression, which we explore in Section 3.
3 Evaluating Setup
3.1 Datasets
To evaluate our approach, we conduct experiments on both open-domain and domain-specific datasets. DBPedia represents an open-domain dataset, encompassing general knowledge across various fields. For domain-specific evaluation, IMDB contains movie-related information while GoodReads focuses on book-related knowledge.
Following Sun et al. (2023), we partition the data into head, torso, and tail portions based on entity frequency, with head containing the top 10% most frequent entities. This dataset selection enables analysis of both domain characteristics and popularity effects on memory patterns.
3.2 Models and Implementation
We conduct experiments using the following LLMs: LLaMA2-13b, LLaMA2-70b, LLaMA3-8b, LLaMA3-70b, LLaMA3.1-8b, Qwen2-1.5b, Qwen2-7b, Qwen2-72b. These models represent different architectural choices and parameter scales, ranging from 1.5B to 70B parameters. To minimize randomness in model outputs, we use greedy decoding with temperature set to 0.0 across all experiments.
3.3 Evaluation Protocol
Given that some models utilize subword tokenization, we employ string comparison to assess whether the model’s output matches the correct answer. Specifically, if any of the top-$k$ tokens share at least three consecutive characters with the ground truth, we classify that token as a match. The value of k in Hits@$k$ correlates with the model’s vocabulary size, particularly important for larger models like LLaMA3 with approximately 128,000 tokens. For questions where the model lacks confidence, we allow it to respond with “unsure” as outlined in our prompt design.
4 Analysis and Results
4.1 Overall Performance
4.1.1 Hits@$k$ performance of different models
A larger model size does not mean a higher Hits@$k$ score
Figure 2 shows the results on dataset DBPedia-head, demonstrating this finding. As the number of parameters increases, LLMs exhibit improved accuracy across a range of tasks, including QA tasks. This is due to the greater representational power of larger models, allowing them to capture more nuanced and complex language patterns. However, among the three datasets used for testing, the Hits@$k$ results for the LLaMA2-13b and LLaMA2-70b models are similar, as are the Hits@$k$ results for the LLaMA3-8b and LLaMA3-70b models. As shown in Figure 3, the rankings of LLMs based on Accuracy and Hits@$k$ differ significantly. This indicates that increasing the model size does not necessarily lead to richer or more comprehensive memory in LLMs.
Figure 2: The Hits@$k$ scores of different large language models on the DBPedia-head dataset.
(a) LLMs sorted by Accuracy
(b) LLMs sorted by Hits@$k$
Figure 3: The ranking of LLMs based on Accuracy and Hits@$k$ on DBPedia-Head.
Newer LLMs have higher Hits@$k$ scores
Our experimental results indicate that newer LLMs exhibit higher Hits@$k$. For instance, the Hits@$k$ of LLaMA3 significantly surpasses that of LLaMA2, regardless of model size. In the head section of the DBPedia dataset, the LLaMA3-70b model achieves a score of 92.1%, the LLaMA3-8b model scores 90.5%, and the LLaMA2-70b model scores 70.5%. This suggests that newer models have a more comprehensive memory of relevant knowledge, likely due to updates in training data. In particular, newer datasets tend to encompass a broader range of information.
4.1.2 Analysis of the influence of $k$ value selection
Figure 4 presents the Hits@$k$ scores for various $k$ values. A lower Hits@1 score suggests that the model struggles to provide the correct answer directly in the QA task. However, as the $k$ value increases, the score improves, indicating that the model retains relevant knowledge. Our experimental results indicate that, when $k=50$, the Hits@$k$ for the head, torso, and tail sections exceeds 80%. Despite the extensive vocabulary of the LLaMA3 model (approximately 128,000 tokens), the correct answer is frequently located within a relatively small number of tokens at the beginning. This suggests that the model has the potential to provide correct answers in most cases, even if an incorrect answer is initially generated. We observe a significant difference between scores at $k=1$ and $k=5$, indicating that utilizing tokens with higher probabilities can yield more reliable answers.
We show the cumulative distribution of the ranks of Hits@$k$ score in the QA task in Figure 5. We observe that the difference in popularity has a smaller impact on Hits@$k$ for the DBPedia dataset compared to IMDB. This suggests that the domain of datasets influences the sensitivity to popularity. Generally, memory performance on open-domain datasets is less sensitive to variations in popularity.
Figure 4: For different values of $k$, We report the Hits@$k$ of LLaMA3-8b on the DBpedia dataset.
(a) DBPedia on LLaMA3-8b
(b) IMDB on LLaMA3-8b
Figure 5: The cumulative distribution of the ranks of Hits@$k$ in the QA task
4.2 Cross-Domain Analysis
4.2.1 Comparison of Open Domain and Specific Domain
As shown in Table 1, The Hits@$k$ results of data sets in different domains are different. DBLP is an open-domain dataset, while IMDB and Goodreads are domain-specific datasets. The experimental results demonstrate that the Hits@$k$ for the open-domain dataset is higher than that for the domain-specific datasets.
We show the cumulative distribution of the ranks of Hits@$k$ score in the QA task in Figure 5. We observe that the difference in popularity has a smaller impact on Hits@$k$ for the DBPedia dataset compared to IMDB. This suggests that the domain of the dataset influences the sensitivity to popularity. Generally, memory performance on open-domain datasets is less sensitive to variations in popularity.
$k=100$ DBPedia IMDB GoodReads
Head Torso Tail Head Torso Tail Head Torso Tail
LLaMA2-13b 70.9 67.0 63.8 44.8 41.7 33.6 36.5 36.5 28.6
LLaMA2-70b 70.5 67.7 64.6 49.4 46.2 35.9 36.1 35.6 31.0
LLaMA3-8b 90.5 88.1 87.1 69.7 66.4 53.5 67.8 68.5 65.6
LLaMA3-70b 92.1 89.2 87.9 56.9 54.1 49.1 44.2 45.4 43.0
LLaMA3.1-8b 89.9 87.5 86.0 69.3 67.0 53.0 67.8 68.3 65.3
Qwen2-1.5b 75.3 71.9 70.8 53.9 48.0 43.4 37.8 38.1 35.3
Qwen2-7b 87.1 85.9 83.9 54.9 52.9 47.2 41.9 43.0 41.7
Qwen2-72b 90.1 88.2 86.8 55.7 52.9 48.4 43.8 44.8 41.7
Table 1: Experimental results (Hits@$k$, $k=100$) for models of varying sizes were obtained by testing different popularity subsets of the head-to-tail dataset.
4.2.2 The impact of domain on knowledge storage
Specific domain datasets are more susceptible to memory loss
Our experimental results show that, compared to open-domain datasets, the Hits@$k$ of specific-domain datasets is lower, indicating that LLMs are more prone to memory loss in specific-domain datasets. This phenomenon may be due to the fact that certain knowledge in specific-domain datasets is not included in the model’s training data.
4.3 Popularity Impact
Popularity impacts the model’s memory storage, though to a lesser extent.
Our experiments indicate that the popularity of datasets influences Hits@$k$. Within the same domain, higher popularity correlates with higher Hits@$k$. However, the difference in Hits@$k$ is smaller than the difference observed when directly calculating the model’s accuracy in QA tasks. This suggests that, beyond the training data, the degree of memory expression significantly impacts the model’s accuracy across datasets with varying popularity. Specifically, in datasets with lower popularity, the model is more likely to retain knowledge related to the questions but may still fail to provide the correct answers.
Popularity exerts a greater influence in specific-domain datasets
Our experiments demonstrate that popularity has a greater impact on Hits@$k$ in specific domain datasets. This suggests that, compared to open-domain datasets, popularity significantly influences memory storage in specific-domain datasets, making it more likely for the model to lack relevant memory in less popular specific-domain datasets.
4.4 Uninformative Response Impact
A noteworthy phenomenon is that, in some cases, the model’s response is uninformative. This includes instances of: 1) repeating specific strings, and 2) outputting empty strings, among others. These types of responses are labeled as “error”. Such responses may arise from anomalies in the model’s generation process, or from a lack of relevant memory.
To reduce the likelihood of the model providing incorrect answers, prompts in QA tasks often include an “unsure” option, allowing the model to respond with “unsure” when uncertain about the correct answer. This approach helps minimize the risk of hallucinations when the model encounters unfamiliar information. Figure 7 shows the distribution ratios of three response types under the LLaMA3-8b model: uninformative, correct, and wrong. We found that uninformative responses have a greater impact on the model’s performance. Our experiments revealed that when some models answered “unsure”, they still retained relevant knowledge in memory. This suggests that the model may respond with “unsure” even when relevant memory exists. We show an example of this situation in Figure 6.
Figure 6: Case study when the LLaMA3-8b model produces uncertain answers. In Question 1 and Question 2, the model’s final answer is “unsure”, but the correct answer, or a subword related to it, appears in the second position of the logits. In Question 3, the model’s final answer is a blank character, which is deemed incorrect. However, the correct answer appears in the token with the third-highest probability.
In summary, we classify the cases mentioned above as uninformative responses, with Figure 6 showing the proportion of such responses across different datasets. We show two different types of uninformative responses in Figure 6, the correct answers appear in the tokens corresponding to the second-highest or third-highest logits. In Questions 1 and 2, the model’s final answer is “unsure,” yet the correct answer, or a related subword, appears as the second most probable token in the logits. In Question 3, the model’s final answer is a null character, which is considered incorrect. However, the correct answer is found in the token with the third-highest probability. This indicates that, while the model possesses relevant memory, it fails to output the correct answer.
As shown in Figure 7, in the DBPedia dataset, experimental results show that more than half of the responses in the Head, Torso, and Tail sub-datasets are uninformative. In the domain-specific IMDB dataset, the high proportion of uninformative responses also significantly impacts the model’s accuracy in QA tasks. This highlights the significant impact of uninformative responses on the final results in both open-domain and domain-specific datasets. Moreover, as dataset popularity decreases, the proportion of uninformative responses increases, which emerges as a key factor contributing to the decline in accuracy in QA tasks.
Our experiments, however, indicate that even uninformative responses may still contain relevant knowledge memory. Fully automating the identification and filtering of incorrect answers is challenging, but identifying and filtering uninformative responses is comparatively straightforward. Since identifying these responses is straightforward, filtering them and extracting the model’s latent knowledge for QA tasks can effectively improve the model’s performance.
Figure 7: We present the distribution of three response types: uninformative, right, and wrong. Additionally, we analyzed data from both open domain and specific domain datasets, reporting the experimental results for the LLaMA3-8b model.
5 Method: SkipUnsure
Algorithm 1 SkipUnsure
1:Inputs: Output’s Token list $L$ (sorted by logits rank)
2:$i=0$
3:while $L[i]$ is uninformative do
4:     Delete $L[i]$
5:     $i=i+1$
6:end while
7:$\text{Prompt}_{new}\leftarrow\text{Prompt}_{old}+L[i]$
8:$\text{Output}_{new}\leftarrow LLM(\text{Prompt}_{new})$
5.1 Methodology
Our analysis in Section 4 demonstrates that LLMs often retain correct knowledge in their parameters even when generating incorrect answers. Particularly, we observe that models frequently output “unsure” or uninformative responses while assigning high probabilities to correct answers in their internal distributions. This finding suggests an opportunity to improve model performance by better utilizing these internal representations.
Motivated by this observation, we propose SkipUnsure, a method to effectively recover stored knowledge from model’s internal representations. Given a question $q$, let $P(t\mid q)$ denote the probability distribution over the model’s vocabulary V, and $T_{k}=\{t_{1},...,t_{k}\}$ be the top-$k$ tokens ranked by their logit scores.
Our method first identifies uninformative tokens through a filtering process. We define a token $t$ as uninformative if it starts with “uns”, represents an empty string, contains less than three characters, or consists only of stop words. This filtering strategy effectively removes low-confidence and meaningless responses while retaining potentially valuable information.
For knowledge recovery, we examine the remaining tokens in $T_{k}$ after filtering out the uninformative set $U$. The final answer $a^{*}$ is selected as the highest-probability token from the valid candidates:
$a^{*}=\arg\max_{t\in T_{k}\setminus U}P(t\mid q)$ (2)
Token $a^{*}$ will be appended to the end of the original prompt and re-entered into the model to generate a new response. This approach allows us to bypass uninformative responses and leverage the model’s stored knowledge that might otherwise be obscured.
The effectiveness of SkipUnsure stems from two key insights from our analysis. First, correct answers often appear in $T_{k}$ with high probability scores even when the model outputs uninformative responses. Second, uninformative tokens can be reliably identified through our filtering criteria. The method requires no model retraining or architectural modifications, making it easy to integrate with existing LLM deployments. Our algorithm is shown in Algorithm 1.
5.2 Experiments result
Following the experimental evaluation settings outlined in the previous article, we conducted evaluations on multiple base models, including those of varying sizes and both recent and older release dates. The prompts used are also consistent with those described in the previous experiments. As shown in Table 3 and Table 3, our experiments demonstrate that our method achieves performance improvements across datasets with varying levels of popularity. The improvement is particularly pronounced in datasets with higher popularity, suggesting that the model has more extensive memory storage in datasets with greater popularity. Using the LLaMA3-8b model, our method improved accuracy on the DBPedia dataset by 3.8% in the head, 3.2% in the torso, and 2.5% in the tail sections. With the LLaMA3-70b model, the improvements were 11.8%, 9.4%, and 6.7% in the respective sections. On the IMDB dataset, with the LLaMA3-70b model, the improvements were 6.3%, 5.4%, and 0.6% in the respective sections. Additionally, our experimental results indicate that uninformative responses impact the model’s memory expression to some extent. By addressing these uninformative responses, the model’s ability to express memory can be enhanced.
DBPedia w/o SkipUnsure SkipUnsure
Head Torso Tail Head Torso Tail
LLaMA2-13b 8.4 4.0 3.2 13.5 $5.1$ 8.9 $4.9$ 6.9 $3.7$
LLaMA2-70b 16.0 12.3 8.1 18.3 $2.3$ 14.0 $1.7$ 9.4 $1.3$
LLaMA3-8b 9.8 7.3 5.1 13.6 $3.8$ 10.5 $3.2$ 7.6 $2.5$
LLaMA3-70b 11.2 8.7 6.0 23.0 $11.8$ 18.1 $9.4$ 12.7 $6.7$
LLaMA3.1-8b 8.1 5.3 3.7 15.6 $7.5$ 10.1 $4.8$ 7.4 $3.7$
Qwen2-1.5b 1.2 0.7 0.6 2.4 $1.2$ 1.5 $0.8$ 1.4 $0.8$
Qwen2-7b 3.9 2.5 1.3 9.3 $5.4$ 7.5 $5.0$ 4.8 $3.5$
Qwen2-72b 17.3 12.1 9.0 20.1 $2.8$ 14.3 $2.2$ 10.2 $1.2$
Table 2: Performance results of SkipUnsure on different base models of dataset DBPedia.
IMDB w/o SkipUnsure SkipUnsure
Head Torso Tail Head Torso Tail
LLaMA2-13b 15.7 11.1 0.0 21.9 $6.2$ 17.6 $6.5$ 2.0 $2.0$
LLaMA2-70b 25.2 23.5 4.3 25.2 23.5 4.5 $0.2$
[…]
LLaMA3-70b 19.1 18.6 4.3 25.4 $6.3$ 24.0 $5.4$ 4.9 $0.6$
[…]
Qwen2-7b 7.4 2.9 0.3 11.7 $4.3$ 4.5 $1.6$ 0.4 $0.1$
Qwen2-72b 19.1 16.2 1.1 20.3 $2.2$ 18.1 $1.9$ 1.2 $0.1$
Table 3: Performance results of SkipUnsure on different base models of dataset IMDB.
6 Related Work
Question-Answering tasks and Hallucination for LLMs
Question-Answering (QA) tasks have become a central application area for LLMs. A key challenge in their adoption and optimization is addressing hallucination, where LLMs generate incorrect or unsupported informationHuang et al. (2023). Currently, there are numerous benchmarks available for evaluating QA tasks on LLMs Berant et al. (2013); Joshi et al. (2017); Dubey et al. (2019); Kwiatkowski et al. (2019); Sciavolino et al. (2021); Mallen et al. (2022); Kumar et al. (2024); Zhong et al. (2023). Sun et al. (2023) proposed datasets partitioned based on popularity. Tonmoy et al. (2024) analyzed the challenges and limitations for hallucination mitigation. Zhang et al. (2023b) analyzed various types of hallucinations in LLMs. Gu et al. (2022) proposed a generic framework and trained a discriminator to evaluate probability of candidate plans for QA tasks. Du et al. (2023) used correlation analysis techniques to quantify and locate the sources of hallucinations, aiming to enhance the reliability of the model. Zhu et al. (2024) evaluated the model’s performance on hallucination problems in real-world scenarios, especially on knowledge-intensive question answering tasks. Waldo and Boussard (2024) analyzed the root causes of hallucinations in large language models and discussed possible directions for improvement.
LLMs as Knowledge Bases
Previous work has proposed that pre-trained language models can be used as knowledge bases Petroni et al. (2019); AlKhamissi et al. (2022). Petroni et al. (2019) introduced the LAMA benchmark, which consists of questions formatted as "fill-in-the-blank" cloze statements. He et al. (2024) explores the potential of LLMs in memorizing exact knowledge in large-scale knowledge bases. Zhong et al. (2023) pointed out the multi-hop knowledge editing problem when using LLMs as knowledge bases. Zheng et al. (2024) investigates the potential of LLMs as knowledge bases, especially in knowledge-intensive tasks. Singhal et al. (2023) demonstrates the potential of LLMs in encoding medical knowledge and answering medical questions. Pan et al. (2024) presents a forward-looking roadmap for the unification of LLMs and Knowledge Graphs (KGs). Yin et al. (2024) offers a fresh perspective and a new method for evaluating large language models, which can help in more accurately understanding and assessing the performance of these models. Hu et al. (2023) evaluates the factual knowledge of LLMs using a benchmark called Pinocchio, which includes 20,000 diverse questions. It finds that while LLMs can implicitly store facts, they often lack accuracy and are unable to update or reason over multiple facts effectively.
7 Conclusion
In this paper, we explore how models express memory in question-answering tasks, defining metrics to assess memory storage and evaluate performance across datasets of varying popularity and domains. Our findings show that the model generally retains relevant knowledge, with memory expression strongly correlating with dataset popularity and domain. We also examine common instances of uninformative responses. Based on these insights, we propose a method to filter out uncertain and irrelevant information, utilizing latent knowledge to improve question answering. Experimental results demonstrate the method’s effectiveness across various models and datasets.